# Student reasoning about quantum mechanics while working with physical experiments

### v1

@@ IV.1 Anwar and Ori
There, Anwar and Ori pointed out that the pattern of coincidence counts increasing and then decreasing was similar to what they had seen with the alignment laser (lines 7-10 and 14). This is a part of the resource Knowledge of classical wave interference that includes knowledge students have about the visual appearance of interference after a laser passes through an interferometer. This resource, however, may also have contributed to the way the students did not initially realize that the experimental results they were observing did not line up with their prediction that single photons would not interfere. The students were still thinking about the alignment laser even as they began working with single photons and therefore did not realize their prediction did not match the experimental results they were seeing until it was pointed out by V.B. (lines 12-20).
@@
When prompted to reason through why they were seeing single-photon interference, Anwar and Ori used the resources Knowledge of classical wave interference and Knowledge of wave functions to make sense of the experimental evidence. In response to V.B.’s question about what is a requirement for interference, Anwar and Ori discussed how it is necessary to have two waves for interference to occur (lines 25-35). When further prompted to think about how they could relate this to the experiment, the students brought up wave functions for the first time (lines 40-44), discussing how a photon’s wave function represents the different probabilities of it traversing each path of the interferometer. The students then went on to explain this by thinking about wave functions mathematically, in particular how wave functions have spatial and temporal components (lines 55-60) and phases (lines 61-63). Putting all of this together allowed Anwar to finally conclude: “So it doesn’t make immediate sense, but wave functions. Wild.”
@@
Anwar and Ori continued to use similar reasoning as they progressed to the next task in the lab guide: blocking one arm of the interferometer when interference was at a minimum for one set of coincidence counts. To explain what they were seeing, they again used the resource Knowledge of wave functions, including that wave functions have phases (lines 6 and 22 in Appendix A.2) and that wave functions can be interpreted as probabilities (lines 12-15, 24, and 48-75). They discussed how, when they were blocking one arm of the interferometer, there was one wave function instead of two, so there was no phase. This removed the interference pattern and thus affected the probabilities they were seeing as counts on the detectors. By the end of this conversation, both students were content to use the idea of wave functions to understand the experimental results they were observing.
@@
Next, Anwar and Ori moved on to looking at the second-order correlation parameter, which provided evidence that the photons were exhibiting particle-like behavior at the same time they were also interfering. Immediately after reading off the second-order correlation parameter from the computer interface, the students again used the resource Knowledge of wave functions, in particular that wave functions represent probabilities (lines 14-23 in Appendix A.3), to reason through why these two types of behavior can exist at the same time. Ori concluded “the wave function is acting like a wave, but it’s really just determining the probability of the particle going into one or the other.” Anwar also mentioned how they were thinking about the spatial dependence of wavefunctions (lines 32-37).
@@
The last part of the experiment involved the students implementing a quantum eraser. After putting half-wave plates in both arms of the interferometer and rotating one of them so the two paths had orthogonal polarizations, Anwar and Ori noticed that the interference had been eliminated. They used the resource Knowledge of classical wave interference, in particular the idea that orthogonal things do not interfere, to understand what they were seeing. The students gestured with their arms to represent the orthogonal polarizations while discussing this with each other. Ori later explained “they’re just not interfering… because, like, the waves are perpendicular to each other.”
@@
Anwar and Ori continued to use the resource of orthogonal things not interfering (a component of Knowledge of classical wave interference) to understand the experimental results after placing a polarizer oriented at in front of detector B’, but not detector B. The students noted that the behavior of the counts in the two detectors differed. Ori described what was happening: “So AB’ is changing a lot. And AB is not. Which suggests that interference is having a big effect going into B’ but not B.” Anwar made sense of this by saying,
@@
After seeing that the interference does indeed reappear in detector B after the insertion of the second polarizer, Ori explained,
@@
Both Anwar and Ori ended the lab session satisfied that they had been able to explain the experimental results they saw.
@@ IV.2 Kiran and Luce
The first clearly identifiable resource Kiran and Luce activated was in the next part of the experiment where they recorded numbers of counts when blocking and unblocking the two arms of the interferometer. In response to V.B.’s question about what effect could be seen in the detectors when the piezo position was adjusted while one arm of the interferometer was blocked, Luce said,
@@
Kiran then tied this idea in with the question asked by stating, “Of course there’s no recombination if we block one of the arms.” Although Kiran and Luce did not articulate a full explanation for single-photon interference at this point, they were activating the resource Knowledge of classical wave interference to understand the changes in the counts they were seeing. They continued to investigate the effects of various actions on the coincidence counts by slowly varying the piezo voltage and blocking one of the arms of the interferometer when at an interference minimum.
Another interesting sense-making moment was when Kiran and Luce looked at the second-order correlation parameter while also seeing interference (see Appendix A.4). There, the students used the resources Knowledge of classical wave interference and Probabilities represent possible quantum outcomes. When asked what was happening at the first beam splitter, Luce pointed to different parts of apparatus, referring to them as either “laser” or “photons.” They were distinguishing between the parts of the apparatus where they were thinking of the photons as having wave-like properties (in the interferometer) and where they were thinking of the photons as having particle-like properties (at the detectors). After acknowledging that there was no laser in the interferometer, yet they were still seeing interference, Luce said, “So somehow the waves are going from waves to single photons here.” To which Kiran responded, “Or could it be that the single photons were waves all along?” Kiran followed this up by discussing probabilities indirectly (“what might happen”) as a way to avoid thinking about splitting individual photons.
@@
They were activating the part of the resource Knowledge of classical wave interference about orthogonal entities not interfering. While discussing how polarizers work, Kiran also activated the resource Probabilities represent possible quantum outcomes:
@@ IV.3 Discussion of student resources
By the end of the experiment, both groups of students were ultimately able to make sense of the fact that they were seeing both particle-like and wave-like behavior of photons at the same time. However, they activated different, yet overlapping, resources to do so. Both groups used their prior knowledge about classical wave interference when first reasoning through the existence of interference and later explaining the quantum eraser. Anwar and Ori spent a large portion of their lab session discussing wave functions, activating both conceptual and mathematical components of this resource including the idea that wave functions represent probabilities of the paths the photons may traverse. Kiran and Luce, on the other hand, never explicitly mentioned wave functions, but instead activated the related resource Probabilities represent possible quantum outcomes. They additionally used the resource Thinking in terms of information when explaining the quantum eraser. The resources activated by both groups at different points in the experiment are shown schematically in Fig. 2.
@@
The resource Knowledge of classical wave interference was used by both groups of students at various stages of the experiment. Kiran and Luce used it throughout, whereas Anwar and Ori used it at the start and end (with discussions of wave functions in the middle). This resource includes a variety of more specific ideas about wave interference and how interferometers work, including ideas such as that waves can constructively and destructively interfere, it is necessary to have two things that are the same to interfere, and orthogonal things do not interfere. These components of the resource (as well as components of the other resources) are listed in Table 1. The students had acquired most of these ideas in previous courses, such as electricity and magnetism. Additionally, students used the idea that interferometers work by splitting and then recombining light, which some of the students may have gained from previous experiences with different types of interferometers. These resources helped students recognize constructive and destructive interference and realize that in order for single photons to interfere, they needed to somehow be divided and then be the same when coming back together.
| Resource | Components |
| Knowledge of classical wave interference |
Waves can constructively and destructively interfere
Need two things for interference to occur Orthogonal things do not interfere Interferometers involve splitting and recombining light Knowledge of what laser interference looks like |
| Knowledge of wave functions |
Wave functions represent probability distributions
Wave functions can constructively and destructively interfere Wave functions have phases Wave functions have spatial components Wave functions have temporal components |
| Probabilities represent possible quantum outcomes | - |
| Thinking in terms of information | - |
@@
The other most frequently activated resource was Knowledge of wave functions, which encompassed both conceptual and mathematical ideas. Anwar and Ori used the ideas that wave functions represent probability distributions, wave functions can constructively and destructively interfere like waves, wave functions have phases, and wave functions can have spatial and temporal components. These resources were used to reason that by changing the path length of one arm of the interferometer, the students were adding a relative phase between two parts of the wave function (although they referred to this as two separate wave functions). This was enabled by the fact that wave functions can change with time and depend on spatial position. The phase affected the probability of the photons going into the two detectors since wave functions can constructively and destructively interfere.
Both students who used the resource Knowledge of wave functions discussed how they had acquired this resource in their quantum mechanics courses. When asked in the post-interview which specific parts of their quantum mechanics courses were most relevant when working through these experiments, Ori said,
@@
Although Kiran and Luce did not explicitly mention wave functions while working through the third experiment, they used the resource Probabilities represent possible quantum outcomes in a similar way. This resource was used to help them understand that even though a single photon cannot be thought of as splitting in half at a beam splitter, the probability of the photon going the different ways can be split. This resource may be a less developed version of the idea that wave functions represent probability distributions. However, it is not clear why Kiran and Luce did not discuss wave functions while working with the experiment, since both of them did mention wave functions when explaining concepts from the experiment in their post-interviews.
The last resource, Thinking in terms of information was used by only Kiran and Luce, and it was an important part of their sense-making about the quantum eraser. In the post-interview, Kiran described this resource as “contextualizing the eraser as a matter of information” since they were “prepared to think about information… as an actual quantifiable thing.” They used this resource to explain how they only saw interference when there was no available information about which path of the interferometer the photons took. This resource may also be a replacement for some of the components of the resource related to wave functions, since in the post-interview Kiran discussed the connection between the concept of information and quantum states:
@@ IV.4 Instructional implications related to student resources
Instructors could utilize the single-photon interferometer experiment after students have learned about classical wave interference and how quantum states are connected to probabilities of outcomes. Both sets of students in our study activated resources related to those ideas in order to understand their experimental results, so it may be useful for instructors to ensure their students have learned these topics beforehand. For instructors incorporating this experiment in a quantum mechanics course, the experiment could be performed after the relevant units. For instructors implementing this experiment in a beyond-first-year lab courses (or any other context where the students may not have previously taken a quantum mechanics course), they could assign preparation activities ahead of time that help students develop these resources.
@@
Instructors could focus more on what students already know than potential misconceptions, or lack of knowledge, while teaching about particle-wave duality of photons. The students we studied did have the resources they needed to be able to make sense of experimental evidence of single-photon interference, even though they did not necessarily phrase their ideas in the same way their instructors might. For example, neither group initially used the word “superposition” and only one emphasized “wave functions,” but both groups still discussed how the photons had some probability of exiting the beam splitter in each of the two directions. Instructors can help students activate and expand on their existing resources that come from both quantum mechanics and other courses.
@@ V Results: Student ideas of what is quantum versus classical
We additionally investigated student ideas related to what is quantum versus classical about the three experiments performed by the students. Not only is understanding the differences between quantum and classical models of light a common learning goal of instructors using these experiments [15], but it is also important to understand what students think is quantum about these experiments since students have been shown to obtain benefits from seeing quantum effects themselves [20]. In this section, we describe the three most prevalent themes we identified in the sequence of think-aloud lab sessions:
@@ V.1 Waves and interference are classical
The first theme we identified encompasses two primary ideas: (1) waves and interference are classical concepts and (2) quantum objects cannot be split. Although these may seem like different ideas, they were often used in conjunction with each other since one line of student reasoning was that waves can split and splitting is a necessary precursor to recombining and therefore interfering. Both of these ideas are similar to already demonstrated student misconceptions of ignoring both the wave nature of photons and the interference of single photons [55].
@@
They started to understand that waves and interference are also a part of the quantum description of light.

### v2

@@ IV.1 Anwar and Ori
There, Anwar and Ori pointed out that the pattern of coincidence counts increasing and then decreasing was similar to what they had seen with the alignment laser (lines 7-10 and 14). They were activating the resource Knowledge of what laser interference looks like since they referenced the visual appearance of varying intensities of the laser after it had passed through the interferometer. This resource, however, may also have contributed to the way the students did not initially realize that the experimental results they were observing did not line up with their prediction that single photons would not interfere. The students were still thinking about the alignment laser even as they began working with single photons and therefore did not realize their prediction did not match the experimental results they were seeing until it was pointed out by V.B. (lines 12-20).
@@
When prompted to reason through the experimental evidence of single-photon interference, Anwar and Ori used various resources in the categories Knowledge of classical wave interference and Knowledge of wave functions. In response to V.B.’s question about what is a requirement for interference, Anwar and Ori activated the resources Need two things for interference to occur and Waves can constructively and destructively interfere (lines 25-35). When further prompted to think about how they could relate this to the experiment, the students brought up wave functions for the first time (lines 40-44), discussing how a photon’s wave function represents the different probabilities of it traversing each path of the interferometer (the resource Wave functions represent probability distributions). The students then went on to explain this by thinking about wave functions mathematically, activating the resources Wave functions have spatial components and Wave functions have temporal components (lines 55-60) as well as Wave functions have phases (lines 61-63). Putting all of this together along with the idea that wave functions can constructively and destructively interfere (lines 64-83), Anwar concluded: “So it doesn’t make immediate sense, but wave functions. Wild.”
@@
Anwar and Ori continued to use similar reasoning as they progressed to the next task in the lab guide: blocking one arm of the interferometer when interference was at a minimum for one set of coincidence counts. To explain what they were seeing, they again used resources related to wave functions, including Wave functions have phases (lines 6 and 22 in Appendix A.2), Wave functions represent probability distributions (lines 12-15, 24, and 48-75), and Wave functions can constructively and destructively interfere (lines 21-38, and 48-68). They discussed how, when they were blocking one arm of the interferometer, there was one wave function instead of two, so there was no phase. This removed the interference pattern and thus affected the probabilities they were seeing as counts on the detectors. By the end of this conversation, both students were content to use the idea of wave functions to understand the experimental results they were observing.
@@
Next, Anwar and Ori moved on to looking at the second-order correlation parameter, which provided evidence that the photons were exhibiting particle-like behavior at the same time they were also interfering. Immediately after reading off the second-order correlation parameter from the computer interface, the students again used their wave function resources, in particular Wave functions represent probability distributions (lines 14-23 in Appendix A.3), to reason through why these two types of behavior can exist at the same time. Ori concluded “the wave function is acting like a wave, but it’s really just determining the probability of the particle going into one or the other.” Anwar also mentioned how they were thinking about the spatial dependence of wave functions (lines 32-37).
@@
The last part of the experiment involved the students implementing a quantum eraser. After putting half-wave plates in both arms of the interferometer and rotating one of them so the two paths had orthogonal polarizations, Anwar and Ori noticed that the interference had been eliminated. They used their knowledge of classical wave interference, in particular the resource Orthogonal things do not interfere, to understand what they were seeing. The students gestured with their arms to represent the orthogonal polarizations while discussing this with each other. Ori later explained “they’re just not interfering… because, like, the waves are perpendicular to each other.”
@@
Anwar and Ori continued to use the resource Orthogonal thing do not interfere to understand the experimental results after placing a polarizer oriented at in front of detector B’, but not detector B. The students noted that the behavior of the counts in the two detectors differed. Ori described what was happening: “So AB’ is changing a lot. And AB is not. Which suggests that interference is having a big effect going into B’ but not B.” Anwar made sense of this by saying,
@@
After seeing that the interference did indeed reappear in detector B after the insertion of the second polarizer, Ori explained,
@@
Here, they also used the resource Things at same place and time can interfere. Both Anwar and Ori ended the lab session satisfied that they had been able to explain the experimental results they saw.
@@ IV.2 Kiran and Luce
The first clearly identifiable resources Kiran and Luce activated were in the next part of the experiment where they recorded numbers of counts when blocking and unblocking the two arms of the interferometer. In response to V.B.’s question about what effect could be seen in the detectors when the piezo position was adjusted while one arm of the interferometer was blocked, Luce said,
@@
Kiran then tied this idea in with the question asked by stating, “Of course there’s no recombination if we block one of the arms.” Although Kiran and Luce did not articulate a full explanation for single-photon interference at this point, they were activating the resources Waves can constructively and destructively interfere and Waves can split and recombine to understand the changes in the counts they were seeing. They continued to investigate the effects of various actions on the coincidence counts by slowly varying the piezo voltage and blocking one of the arms of the interferometer when at an interference minimum.
Another interesting sense-making moment was when Kiran and Luce looked at the second-order correlation parameter while also seeing interference (see Appendix A.4). There, the students again used the resources Waves can constructively and destructively interfere and Waves can split and recombine, while also activating the resources Knowledge of what laser interference looks like and Quantum outcomes are probabilistic. When asked what was happening at the first beam splitter, Luce pointed to different parts of apparatus, referring to them as either “laser” or “photons.” They were distinguishing between the parts of the apparatus where they were thinking of the photons as having wave-like properties (in the interferometer) and where they were thinking of the photons as having particle-like properties (at the detectors). After acknowledging that there was no laser in the interferometer, yet they were still seeing interference, Luce said, “So somehow the waves are going from waves to single photons here.” To which Kiran responded, “Or could it be that the single photons were waves all along?” Kiran followed this up by discussing probabilities indirectly (“what might happen”) as a way to avoid thinking about splitting individual photons.
@@
They were activating the resource Orthogonal things do not interfere. While discussing how polarizers work, Kiran also activated the resource Quantum outcomes are probabilistic:
@@ IV.3 Discussion of student resources
By the end of the experiment, both groups of students were ultimately able to make sense of the fact that they were seeing both particle-like and wave-like behavior of photons at the same time. However, they activated different, yet overlapping, resources to do so. Both groups used their prior knowledge about classical wave interference when first reasoning through the existence of interference and later explaining the quantum eraser. Anwar and Ori spent a large portion of their lab session discussing wave functions, activating both conceptual and mathematical resources including the idea that wave functions represent probabilities of the paths the photons may traverse (the resourceWave functions represent probability distributions). Kiran and Luce, on the other hand, never explicitly mentioned wave functions, but instead activated the related resource Quantum outcomes are probabilistic. They additionally used the resource Thinking in terms of information when explaining the quantum eraser. The broad categories of resources (or resources for the two that did not fit into categories) that both groups of students activated at different points in the experiment are shown schematically in Fig. 2 to provide a visual overview of the two groups’ reasoning. The specific resources activated by each pair of students are indicated in Table 2.
| Varied piezo, saw counts changing | Blocked arms of interferometer | Second-order correlation measurement | Quantum eraser | ||
| Anwar and Ori | |||||
| Waves can constructively and destructively interfere | ✓ | ||||
| Need two things for interference to occur | ✓ | ||||
| Things at same place and time can interfere | ✓ | ||||
| Orthogonal things do not interfere | ✓ | ||||
| Waves can split and recombine | |||||
| Knowledge of what laser interference looks like | ✓ | ||||
| Wave functions represent probability distributions | ✓ | ✓ | ✓ | ||
| Wave functions can constructively and destructively interfere | ✓ | ✓ | |||
| Wave functions have phases | ✓ | ✓ | |||
| Wave functions have spatial components | ✓ | ✓ | |||
| Wave functions have temporal components | ✓ | ||||
| Quantum outcomes are probabilistic | |||||
| Thinking in terms of information | |||||
| Kiran and Luce | |||||
| Waves can constructively and destructively interfere | ✓ | ✓ | |||
| Need two things for interference to occur | |||||
| Things at same place and time can interfere | |||||
| Orthogonal things do not interfere | ✓ | ||||
| Waves can split and recombine | ✓ | ✓ | |||
| Knowledge of what laser interference looks like | ✓ | ||||
| Wave functions represent probability distributions | |||||
| Wave functions can constructively and destructively interfere | |||||
| Wave functions have phases | |||||
| Wave functions have spatial components | |||||
| Wave functions have temporal components | |||||
| Quantum outcomes are probabilistic | ✓ | ||||
| Thinking in terms of information | ✓ | ||||
@@
The resources in the category Knowledge of classical wave interference were used by both groups of students at various stages of the experiment. Kiran and Luce used these resources throughout, whereas Anwar and Ori used them at the start and end, with discussions of wave functions in the middle. Both groups used the resources Waves can constructively and destructively interfere and Knowledge of what laser interference looks like to understand that the single photons were interfering, and they both also used the resource Orthogonal things do not interfere while investigating the quantum eraser. The students had acquired most of these ideas in previous courses, such as electricity and magnetism. The students may have additionally gained the resources related to laser interference and the splitting and recombining of light from previous experiences with different types of interferometers in lab classes or research experiences. This category of resources helped students recognize constructive and destructive interference and realize that in order for single photons to interfere, they needed to somehow be divided and then be the same when coming back together.
@@
Resources in the category Knowledge of wave functions, which encompasses ideas related to both conceptual and mathematical understanding of wave functions, were also activated frequently by Anwar and Ori. They activated the resources Wave functions represent probability distributions, Wave functions can constructively and destructively interfere, Wave functions have phases, Wave functions have spatial components and Wave functions have temporal components. These resources were used to reason that by changing the path length of one arm of the interferometer, the students were adding a relative phase between two parts of the wave function (although they referred to this as two separate wave functions). This was enabled by the fact that wave functions can change with time and depend on spatial position. The phase affected the probability of the photons going into the two detectors since wave functions can constructively and destructively interfere.
Both students who used the resources in the category Knowledge of wave functions discussed how they had acquired this resource in their quantum mechanics courses. When asked in the post-interview which specific parts of their quantum mechanics courses were most relevant when working through these experiments, Ori said,
@@
Although Kiran and Luce did not explicitly mention wave functions while working through the third experiment, they used the resource Quantum outcomes are probabilistic in a similar way. This resource was used to help them understand that even though a single photon cannot be thought of as splitting in half at a beam splitter, the probability of the photon going the different ways can be split. This resource may be a less developed version of the resource Wave functions represent probability distributions. However, it is not clear why Kiran and Luce did not discuss wave functions while working with the experiment, since both of them did mention wave functions when explaining concepts from the experiment in their post-interviews.
The last resource, Thinking in terms of information was used by only Kiran and Luce, and it was an important part of their sense-making about the quantum eraser. In the post-interview, Kiran described this resource as “contextualizing the eraser as a matter of information” since they were “prepared to think about information… as an actual quantifiable thing.” They used this resource to explain how they only saw interference when there was no available information about which path of the interferometer the photons took. This resource may also be a replacement for some of the resources related to wave functions, since in the post-interview Kiran discussed the connection between the concept of information and quantum states:
@@ IV.4 Instructional implications related to student resources
Instructors could consider utilizing the single-photon interferometer experiment after students have learned about classical wave interference and how quantum states are connected to probabilities of outcomes. The focus of the single-photon interferometer is not on classical wave interference, yet understanding that two waves can constructively and destructively interfere and what that looks like may be necessary for students to understand that they are seeing interference with single photons. Additionally, it may be helpful for students to have at least some idea that measurement outcomes in quantum mechanics are probabilistic. Both sets of students in our study activated resources related to these ideas in order to understand their experimental results, so instructors may want to ensure their students have learned these topics before working with the experiment.
The decision of when to incorporate an experiment into a course is important as it affects what conclusions the students can come to while working with the experiment and what learning they may be primed for afterward. From our data, we cannot definitively say which concepts the students should know ahead of time versus which ones they could best learn in the context of these experimental observations, but students may engage with the ideas more productively when the experiments occur after or in conjunction with units on topics such as wave functions, superpositions, and measurement probabilities (if included in a quantum mechanics course). For instructors implementing this experiment in a BFY lab course (or any other context where the students may not have previously taken a quantum mechanics course), they could assign activities that help students develop these resources alongside the experiment.
@@
Even for potentially unintuitive topics like particle-wave duality, instructors could help students activate their own resources, whether or not they are fully-developed and canonically correct. The students we studied did have the resources they needed to be able to make sense of experimental evidence of single-photon interference, even though they did not necessarily phrase their ideas in the same way their instructors might. For example, neither group initially used the word “superposition” and only one emphasized “wave functions,” but both groups still discussed how the photons had some probability of exiting the beam splitter in each of the two directions. Connections between the specific resources activated in the different parts of the experiment (see Table 2) may also help instructors decide which activities to include as part of their lab guides and when they may be able to help students activate and expand on their existing resources.
@@ V Results: Student ideas of what is quantum versus classical
We additionally investigated student ideas related to what is quantum versus classical about the three experiments, another type of student reasoning about which instructors care. Not only is understanding the differences between quantum and classical models of light a common learning goal of instructors using these experiments [15], but it is also important to understand what students think is quantum about these experiments since students have been shown to obtain benefits from seeing quantum effects themselves [20]. In this section, we describe the three most prevalent themes we identified in the sequence of think-aloud lab sessions:
@@ V.1 Waves and interference are classical
The first theme we identified encompasses two primary ideas: (1) waves and interference are classical concepts and (2) quantum objects cannot be split. Although these may seem like different ideas, they were often used in conjunction with each other since one line of student reasoning was that waves can split and splitting is a necessary precursor to recombining and therefore interfering. Both of these ideas are similar to already demonstrated student difficulties of ignoring both the wave nature of photons and the interference of single photons [45].
@@
They started to understand that wave behavior and interference are also a part of the quantum description of light.